How much of the solutions do I need to use when one container has 40% and another has 100% and 50% is needed for one gallon?I came up with a huge number.
Answer by stanbon(75887)

How much of the solutions do I need to use when one container has 40% and another has 100% and 50% is needed for one gallon?
-------------
Equation:
0.40x + 1(1-x) = 0.50(1)
Multiply thru by 100 to get:
40x + 100(1-x) = 50
40x + 100 -100x = 50
-60x = -50
x = 5/6 gal. (amount of 40% solution needed in the mixture)
---
1-x = 1-(5/6) = 1/6 gal. (amount of 100% solution needed in the mixture)
===========================================================================
Cheers,
Stan H.